# Volume-Preserving Deformation of Honeycomb Wire Media Enables Broad Plasma Frequency Tunability

### v2

@@ Abstract
We demonstrate tunability of the plasma frequency in a wire medium by mechanically deforming a lattice of parallel metallic wires arranged at the nodes of a honeycomb structure. Numerical simulations predict up to 78% tunability and a proof-of-concept experiment confirms 64%.
@@
A wire medium [17, 18] refers to a structure composed of parallel metallic wires periodically arranged in a perpendicular plane. This wire metamaterial prevents wave propagation below a characteristic frequency known as the plasma frequency. This frequency depends on the periodicity of the material, i.e. on the geometry of the unit cell [19], on the number of wires and their arrangement within the unit cell [20, 13]. The plasma frequency determines the dispersion of the wire medium [19, 8] and sets the epsilon-near-zero condition [21, 5], which can be tailored to enhance antenna performance [22, 23].
@@
Figure 1: (a) Geometry of a wire metamaterial with a honeycomb lattice of period $a$. A hexagonal unit cell (solid black line) contains parallel wires of radius $r_{0}$. $R$ is the distance between the unit cell center and wires within a unit cell. (b) Breathing deformation of the structure visualization. (c) Numerical calculation of the metamaterial plasma frequency during the breathing deformation. Geometry parameter $R$ changes from $2r_{0}$ to $a/2-r_{0}$ ($a=3$ cm, $r_{0}=0.5$ $m$). Insets in the plot show field distributions of the plasma mode for each highlighted $R/a$ value.
@@
To demonstrate a tunability provided by the breathing deformation we performed a numerical simulation in COMSOL [25] for an infinite wire medium (applying periodic boundary conditions to a hexagonal unit cell) with a period $a=3$ cm composed of infinite and perfectly electric conducting (PEC) wires of radii $r_{0}=0.5$ mm. The plot in Fig. 1(c) shows a dependence of the plasma frequency on a deformation parameter $R$ changing from the minimal possible size equal to $2r_{0}$, where wires touch near the unit cell center, to the maximal $a/2-r_{0}$, where wires from neighboring cells touch near the unit cell boundary.
@@
Figure 3: (a) Resonance frequency, (b) form factor, and (c) quality factor of the fundamental TM mode in wire medium cavities as a function of the geometry parameter $R$, which varies from $2r_{0}$ to $a/2-r_{0}$. Introducing a quarter-wavelength air gap between the wire medium and the cavity walls causes the cavity to behave similarly to an infinite wire medium structure, with minimal impact on the quality factor. Points $\mathrm{P_{1}}$ and $\mathrm{P_{2}}$ correspond to the $R/a$ values used in the experimental resonators described in the text.
@@
The next step of our study was to implement numerical simulations of a finite resonator to assess whether the observed tunability could be realized experimentally. We simulated a hexagonal cavity filled with the metamaterial, as shown in Fig. 2. The hexagonal cavity shape was chosen to achieve optimal termination of the wire medium, enclosing nearly an integer number of hexagonal unit cells. A typical configuration for wire medium cavities involves enclosing the wire medium with metallic walls placed at the edges of the peripheral unit cells. This arrangement supports a fundamental TM mode with a uniform longitudinal electric field and a sinusoidal transverse field profile, and its resonance frequency lies above the plasma frequency [15]. In contrast, an optimized configuration introduces an air gap of quarter-wavelength thickness, $\Delta=\lambda_{p}/4$, between the peripheral unit cells and the cavity walls. This adjustment yields a uniform field distribution for the fundamental TM mode and ensures that its resonance frequency coincides with the plasma frequency [27]. In this context, TM modes are defined as those in which the electric field is aligned with the wire axes.
@@
Figure 3(a) shows the resonant frequency of the fundamental TM mode in wire medium cavities with an air gap of $\Delta=0$ (regular cavity) and $\Delta=\lambda_{p}/4$ (optimized cavity). The results for the cavities indicate that the tuning percentage remains similar to the infinite wire medium, reaching values above $70\%$ (see Fig. 3(a)). However, the optimized cavity resonates exactly at the plasma frequency, providing a direct method for measuring the plasma frequency of the wire medium.
@@
Although the primary goal of this work is to demonstrate a tunable plasma frequency in a wire medium, the wire medium cavities used for experimental validation are also relevant to axion haloscope applications. Simulations indicate that the form factor and quality factor (see Figs. 3(b,c)) of the regular and optimized cavity designs fall within acceptable ranges for such applications [28], suggesting that the proposed geometry could serve as a viable platform for future dark matter detection experiments.
(a)(b)(c)(d)(e)(f)(g)(h)
@@
Figure 4: Wire medium resonators exhibiting (a–d) low and (e–h) high plasma frequencies. (a,e) Photographs of the experimental resonators; (b,f) simulated field mode profiles; (c,g) simulated $S_{21}$ parameter; (d,h) experimental $S_{21}$ parameter. The plasma frequency corresponds to the fundamental resonance frequency. The geometrical parameters of the resonators are listed in Table 2.
@@
We fabricated two optimized hexagonal cavities (see Figs. 4(a,e)). Each cavity was filled with a wire medium based on the breathing honeycomb configuration (see Figs. 1(a) and 2), but with different values of the geometric parameter $R$. To ensure that the plasma frequency coincided with the fundamental TM mode resonance of the cavities, we optimized the spacing $\Delta$ between the peripheral unit cells and the cavity walls ($\Delta=\lambda_{p}/4$). The geometric parameters of the experimental resonators are summarized in Table 2.
@@
We built the experimental resonators using 1 mm-thick double-sided printed circuit boards (PCBs) for the cavity walls. To ensure electrical contact between the wires and the top and bottom plates, we used 1 mm-thick single-sided PCBs with the conductive side facing outward. The top and bottom plates were drilled to accommodate both the wires and the SMA ports. Each resonator consisted of 19 hexagonal unit cells with a lattice constant $a=3$ cm, enclosed by hexagonal cavity walls. A total of 114 wires with radius $r_{0}=0.5$ mm were soldered into the structures, as shown in Figs. 4(a,e).
We soldered a 3-mm-long SMA connector to the top and bottom PCBs for signal feeding, as shown in Figs. 4(a,e). The SMA ports function as monopole antennas, exciting TM modes within the cavities. We measured the transmission parameter $S_{21}$ using a planar vector network analyzer (VNA, model S5085). Figure 2 shows the exact location of the port in a top view of the wire medium resonators.
@@
Figures 4(c,g) and 4(d,h) show the $S_{21}$ parameter for the resonators obtained from simulations and experiments, respectively. We focus in particular on the fundamental TM-mode resonance, which corresponds to the plasma frequency in our optimized cavities. This identification is further supported by the uniform field distribution of the $\text{TM}_{000}$ mode observed in the numerical simulations shown in Figs. 4(b,f). Resonator I exhibits a plasma frequency of 3.745 GHz in the simulation and 3.750 GHz in the experiment. Resonator II exhibits a plasma frequency of 7.317 GHz in the simulation and 7.285 GHz in the experiment. The relative error in the plasma frequency between simulations and experiments is approximately 0.1% for Resonator I and 0.4% for Resonator II.

### v3

@@ Abstract
We demonstrate significant tunability of the plasma frequency in a wire medium by mechanically deforming a lattice of parallel metallic wires arranged at the nodes of a honeycomb structure. Numerical simulations predict up to 78% tunability and a proof-of-concept experiment confirms 64%, surpassing previously reported values for tunable wire media.
@@
A wire medium [13, 14, 15] refers to a structure composed of parallel metallic wires periodically arranged in a perpendicular plane. This wire metamaterial prevents wave propagation below a characteristic frequency known as the plasma frequency. This frequency depends on the periodicity of the material, i.e. on the geometry of the unit cell [16], on the number of wires and their arrangement within the unit cell [17, 18]. The plasma frequency determines the dispersion of the wire medium [16, 8] and sets the epsilon-near-zero condition [19, 5], which can be tailored to enhance antenna performance [20, 21].
@@
Figure 1: (a) Geometry of a wire metamaterial with a honeycomb lattice of period $a$. A hexagonal unit cell (solid black line) contains parallel wires of radius $r_{0}$. $R$ is the distance between the unit cell center and wires within a unit cell. (b) Breathing deformation of the structure visualization. (c) Numerical calculation of the metamaterial plasma frequency during the breathing deformation. Geometry parameter $R$ changes from $2r_{0}$ to $a/2-r_{0}$ ($a=3$ cm, $r_{0}=0.5$ $m$) in steps of $10^{-3}a$ (451 data points)
. Insets in the plot show field distributions of the plasma mode for each highlighted $R/a$ value.
@@
To demonstrate a tunability provided by the breathing deformation we performed a numerical simulation in COMSOL [26] for an infinite wire medium (applying periodic boundary conditions to a hexagonal unit cell and calculating the lowest eigenmode in the $\Gamma$-point) with a period $a=3$ cm composed of infinite and perfectly electric conducting (PEC) wires of radii $r_{0}=0.5$ mm. The plot in Fig. 1(c) shows a dependence of the plasma frequency on a deformation parameter $R$ changing from the minimal possible size equal to $2r_{0}$, where wires touch near the unit cell center, to the maximal $a/2-r_{0}$, where wires from neighboring cells touch near the unit cell boundary, with a step equal to $10^{-3}a$.
@@
Figure 3: (a) Resonance frequency, (b) form factor, and (c) quality factor of the fundamental TM mode in wire medium (WM) cavities as a function of the geometry parameter $R$, which varies from $2r_{0}$ to $a/2-r_{0}$. Introducing a quarter-wavelength air gap between the wire medium and the cavity walls causes the cavity to behave similarly to an infinite wire medium structure, with minimal impact on the quality factor. Points $\mathrm{P_{1}}$ and $\mathrm{P_{2}}$ correspond to the $R/a$ values used in the experimental resonators described in the text.
@@
The next step of our study was to implement numerical simulations of a finite resonator to assess whether the observed tunability could be realized experimentally. We simulated a hexagonal cavity filled with the metamaterial, as shown in Fig. 2. The hexagonal cavity shape was chosen to achieve optimal termination of the wire medium, enclosing nearly an integer number of hexagonal unit cells. A typical configuration for wire medium cavities involves enclosing the wire medium with metallic walls placed at the edges of the peripheral unit cells. This arrangement supports a fundamental TM mode with a uniform longitudinal electric field and a sinusoidal transverse field profile, and its resonance frequency lies above the plasma frequency [23]. In contrast, an optimized configuration introduces an air gap of quarter-wavelength thickness, $\Delta=\lambda_{p}/4$, between the peripheral unit cells and the cavity walls. This adjustment results in a wire medium bounded by an effective perfect magnetic conductor (PMC) boundary condition [28], which produces a uniform field distribution for the fundamental TM mode and ensures that its resonance frequency coincides with the plasma frequency [28]. In this context, TM modes are defined as those in which the electric field is aligned with the wire axes.
@@
Figure 3(a) shows the resonant frequency of the fundamental TM mode in wire medium cavities with an air gap of $\Delta=0$ (regular cavity) and $\Delta=\lambda_{p}/4$ (optimized cavity). The results for the cavities indicate that the tuning percentage remains similar to the infinite wire medium, reaching values above $70\%$ (see Fig. 3(a)). Notably, the optimized cavity resonates exactly at the plasma frequency, providing a direct method for measuring the plasma frequency of the wire medium. It is worth mentioning that while a regular cavity approaches the plasma frequency as the number of cells inside the cavity increases [23], the optimized cavity reproduces the resonance of the infinite wire medium with only a small number of unit cells, as illustrated in Fig. 3(a).
@@
Although the primary goal of this work is to demonstrate a tunable plasma frequency in a wire medium, the wire medium cavities used for experimental validation are also relevant to axion haloscope applications. Simulations indicate that the form factor and quality factor (see Figs. 3(b,c)) of both the regular and optimized cavity designs fall within acceptable ranges for such applications [29], suggesting that the proposed geometry could serve as a viable platform for future dark matter detection experiments. We obtained these values in COMSOL [26] by modeling the wires and cavity walls with an impedance boundary condition corresponding to copper ($\sigma=5.7\times 10^{7}~\mathrm{S/m}$) when computing the quality factor [23], and by using Ref. [28, Eq. (13)] to calculate the form factor.
(a)(b)(c)(d)(e)(f)
@@
Figure 4: Wire medium resonators exhibiting (a–c) low and (d–f) high plasma frequencies. (a,d) Photographs of the experimental resonators; (b,e) simulated field mode profiles; (c,f) simulated and experimental $S_{21}$ parameter. The plasma frequency corresponds to the fundamental resonance frequency. The geometrical parameters of the resonators are listed in Table 2.
@@
We fabricated two optimized hexagonal cavities (see Figs. 4(a,d)). Each cavity was filled with a wire medium based on the breathing honeycomb configuration (see Figs. 1(a) and 2), but with different values of the geometric parameter $R$. To ensure that the plasma frequency coincided with the fundamental TM mode resonance of the cavities, we optimized the spacing $\Delta$ between the peripheral unit cells and the cavity walls ($\Delta=\lambda_{p}/4$). We obtained the plasma wavelength $\lambda_{p}$ by simulating an infinite wire medium, as shown in Fig. 1 and Table 1 (points II and IV). The geometric parameters of the experimental resonators are summarized in Table 2.
@@
We built the experimental resonators using 1 mm-thick double-sided printed circuit boards (PCBs) for the cavity walls. To ensure electrical contact between the wires and the top and bottom plates, we used 1 mm-thick single-sided PCBs with the conductive side facing outward. The top and bottom plates were drilled to accommodate both the wires and the SubMiniature version A (SMA) connector ports. Each resonator consisted of 19 hexagonal unit cells with a lattice constant $a=3$ cm, enclosed by hexagonal cavity walls. A total of 114 wires with radius $r_{0}=0.5$ mm were soldered into the structures, as shown in Figs. 4(a,d).
We soldered a 3-mm-long SMA connector to the top and bottom PCBs for signal feeding, as shown in Figs. 4(a,d). The SMA ports function as monopole antennas, exciting TM modes within the cavities [23], such that each peak in the $S_{21}$ spectrum corresponds to a resonance mode. We measured the transmission parameter $S_{21}$ using a planar vector network analyzer (VNA, model S5085). Figure 2 shows the exact location of the port in a top view of the wire medium resonators.
@@
Figures 4(c,f) show the $S_{21}$ parameter for the resonators obtained from simulations and experiments. We focus in particular on the fundamental TM-mode resonance, which corresponds to the plasma frequency in our optimized cavities. This identification is further supported by the uniform field distribution of the $\text{TM}_{000}$ mode observed in the numerical simulations shown in Figs. 4(b,e). Resonator I exhibits a plasma frequency of 3.745 GHz in the simulation and 3.750 GHz in the experiment. Resonator II exhibits a plasma frequency of 7.317 GHz in the simulation and 7.285 GHz in the experiment. The relative error in the plasma frequency between simulations and experiments is approximately 0.1% for Resonator I and 0.4% for Resonator II.
@@
Acknowledgements.
The numerical part of this study was carried out with support from the state assignment No. FSER-2024–0041 within the framework of the national project “Science and Universities.” The experimental part was funded by the Russian Science Foundation, grant No. 25-12-00261 (https://rscf.ru/project/25-12-00261/).
Data availability
The data that support the findings of this article are openly available [36].